### v1

# Widespread disruption of resonant chains during protoplanetary disk dispersal

Abstract
We describe the evolution of low mass planets in a dispersing protoplanetary disk around a Solar mass star. The disk model is based on the results of Yu et al. (2023), which describes a region of the inner disk where the direction of the migration torque is outwards due to the diffusion of the stellar magnetic field into the disk and the resultant gradual increase in surface density outwards. We demonstrate that the magnetospheric rebound phase in such a disk leads to diverging orbits for double and triple planet systems, and the disruption of a high fraction of the initial resonant chains.
We present simulations of three planet systems with masses based on the observed triple planet systems observed by the Kepler satellite within the context of this model. The final distribution of nearest neighbour period ratios provides an excellent fit to the observations, provided that the initial resonant configurations are compact. The occurrence rate of planets as a function of orbital period also provides a good match to the observations, for final orbital periods $P<20$ days.
These results suggest that the period and period ratio distributions of low mass planets are primarily set in place during the disk dispersal epoch, and may not require significant dynamical evolution thereafter.
1 Introduction
The Kepler satellite mission (Borucki et al., 2010) has revealed that $>50\%$ of Sun-like stars possess planetary systems containing at least one and, frequently, several planets with orbital periods $<100$ days (Batalha et al., 2013; Coughlin et al., 2016; Lissauer et al., 2023). The majority of these planets have masses that lie between those of Earth and Saturn, suggesting a commonality with either rocky planets (hereafter Super-Earths) or planets with a substantial gaseous or water content by mass (hereafter Sub-Neptunes). Despite their widespread occurrence, the origins of these planets is still a matter of lively debate.
One school of thought suggests that such planets, especially those with volatile-rich envelopes, formed further from the star and migrated inwards due to interactions with the gaseous protoplanetary disk. The migration timescales associated with low mass planets are quite short, but gravitational interactions between multiple migrating planets can slow the evolution sufficiently to avoid losing the planets into the central star (Terquem & Papaloizou, 2007; Ida & Lin, 2008, 2010; Hellary & Nelson, 2012; Cossou et al., 2014; Coleman & Nelson, 2014; Izidoro et al., 2017). The high frequency of multiple planet systems also aligns nicely with this model. However, such models predict a high occurrence rate of mean motion resonances in such systems, which is not matched by observations (Lissauer et al., 2011; Fabrycky et al., 2014). An alternative class of models suggests that the planets assembled in situ, with little radial migration (Hansen & Murray, 2012; Chiang & Laughlin, 2013; Boley et al., 2014; Chatterjee & Tan, 2015). These models can better match the observed spacing of planets in multiple systems (Hansen & Murray, 2013), but require that the protoplanetary disks be either unusually massive or allow for significant inward migration of solid material as small bodies, prior to assembly.
For models in which the planets migrate within the protoplanetary disk, a critical element is the process that determines the halting of migration. In particular, the structure of the inner edge of the disk is expected to be regulated by the interaction with the stellar magnetic field (Romanova & Lovelace, 2006; Romanova et al., 2019) and the change in the density structure reduces or reverses the torque driving the planets inwards (Masset et al., 2006; Tsang, 2011; Miranda & Lai, 2018). For the purposes of modelling migration, this is often modelled as a simple truncation, resulting from the balance of the ram pressure of the disk flow with the magnetic pressure of the magnetic field (Terquem & Papaloizou, 2007; Chang et al., 2010; Izidoro et al., 2017; Brasser et al., 2018). However, Yu et al. (2023) have shown that a more accurate model – accounting for the diffusion of the magnetic field into the disk – results in an inner transition region, where the torque on the planet is not zero but changes sign, driving outwards migration. Thus, migrating planets reverse the course of their migration as they cross this transition and therefore pile up not at the edge of the magnetosperic cavity, but at a more distant location. As the accretion rate drops, the torque reversal location moves outwards, causing the planet to move outwards again, coming to rest at the point where the disk density has dropped to the point that the migration ‘freezes out’.
In Yu et al. (2023) we demonstrated how this model determines the final orbital periods in the case of a single planet. In this paper, we wish to investigate the consequences of this model for multiple planet systems. An important feature of the multiple planet case is that it can result in planetary pairs breaking out of resonance. Liu et al. (2017) have shown that, as the accretion rate drops, the magnetospheric cavity expands and can drive outward migration of a planet trapped at the inner edge. They use a one-sided torque operating at the disk edge, based on the results of the studies above. The fact that outward torques operate only on one planet limits the degree to which this effect operates to disrupt a resonance chain, and the resulting simulations still leave too many systems in resonant configurations (Liu & Ormel, 2017). As we will show, the presence of a non-zero torque interior to the location of the torque reversal results in multiple planets migrating outwards and provides a more extensive dynamical perturbation to the system.
In this paper we present a calculation of the resonant dynamics of two and three planet systems, evolving under the influence of a torque model adapted from Yu et al. (2023). The dynamical model, including both resonant interactions and torques from the protoplanetary disk, is described in § 2. In § 3 we adapt the semi-analytic treatment of the resonant dynamics to take into account the torque reversals. § 4 presents the application of our model to two planet systems and describes the character of the evolution over the range of planetary masses. In § 5 we show how the introduction of a third planet influences the dynamics and present a comparison between model results and the Kepler observational sample. The implications of our results are discussed in § 6.
2 Dynamical Model
We wish to model the resonant dynamics of two and three planet systems, under the influence of the most important resonant interactions. We choose to focus on the first order resonances $(q+1)/q$ where $q=1$ and $q=2$. We will also include the effects of torques from the protoplanetary disk on both semi-major axis evolution and planetary eccentricity damping. This is done by including, in the equations, a linear decay term characterised by timescale, as has been done by several prior authors – (Terquem & Papaloizou, 2019; Choksi & Chiang, 2020). One difference in our case is that our disk model is based on the model of Yu et al. (2023), which features a reversal of the torque direction in the inner disk, due to the penetration of the stellar magnetic field into the disk. As a result, the sign of the torque can reverse, so that we must allow for the timescale for semi-major axis evolution ($\tau_{a}$) to become negative. The timescale for eccentricity damping is always positive. The equations featuring three planets and two resonances are sufficiently lengthy that we defer their explicit expression to appendix A.
2.1 Torque Model
The dynamical evolution of the system is driven by the torques exerted by the protoplanetary disks. We determine a functional form for the torque based on a fit to the numerical disk simulations discussed in Yu et al. (2023). For the inward migration, the characteristic timescale for semi-major axis evolution is
$\tau_{a}$ $=$ $0.36{\rm Myr}\left(\frac{a}{0.1AU}\right)^{1.38}\left(\frac{M_{p}}{10M_{\oplus}}\right)^{-1}$ (1)
$\left(\frac{\dot{M}}{1.8\times 10^{-9}M_{\odot}/yr}\right)^{-0.85}$
The time dependance of $\tau_{a}$ is driven by the evolution of the accretion rate, which is
$\dot{M}=\frac{2.5\times 10^{-8}M_{\odot}yr^{-1}exp(-t/0.63Myr)}{1+exp((t-1.7Myr)/0.15Myr)}.$ (2)
This is chosen to represent a disk that decays exponentially as it accretes onto the star but then cuts off more rapidly once photoevaporation losses detach the inner disk from mass resupply by the outer disk (Alexander et al., 2014).
Once a planet has crossed interior to the torque reversal line $a_{R}$, the migration is now directed outwards. We describe this using a negative timescale. The functional form is again determined by fitting to the simulations of Yu et al. (2023), and is expressed as a function of position relative to $a_{R}$, $z=\ln a_{1}-a_{R}$. This amounts to
$\tau_{a}$ $=$ $6.66\times 10^{-18}{\rm Myr}\left(\frac{0.1AU}{a}\right)^{1/2}\left(\frac{10M_{\oplus}}{M_{p}}\right)$
$\left(\frac{\dot{M}}{1.8\times 10^{-9}M_{\odot}/yr}\right)^{-1.8}e^{-(120.04+177.76z+65.5z^{2}+8.07z^{3})}$
These quantities are furthermore linked by the fact that the transition radius $a_{R}$ is, itself, a function of $\dot{M}$ and thus of time, with
$a_{R}=0.062AU\left[1+0.229\ln\left(\frac{\dot{M}}{2.5\times 10^{-8}M_{\odot}/yr}\right)\right]^{-1}$ (4)
in the case of our default model. Indeed, if the migration time inward or outward is shorter than the timescale on which $a_{R}$ is evolving, then the planet will track the torque reversal location, making the relevant characteristic timescale $\tau_{R}=dt/d\ln a_{R}$.
For example, if the outer planet of the pair reaches the torque reversal location, then the pair of planets will both be migrating outwards. The question of whether they converge or diverge will rest on whether the outward migration of the inner planet is faster or slower than the rate at which $a_{R}$ recedes. Two examples of this case are shown in Figure 1. The solid curve shows the characteristic timescale associated with the outward movement of the torque reversal location as $\dot{M}$ decreases. We also show two cases of $\tau_{a}$ interior to this. The dotted line shows the rate of outward migration for a $6M_{\oplus}$ planet that is the interior member of a pair locked in the 2:1 mean motion resonance, in which the outer member is located at $a_{R}$. We see that this increases with time, while the solid curve decreases with time. Thus, initially, the inner planet is moving outwards faster, and remains locked in resonance because the pair is converging. However, when the two curves cross, the inner planet can no longer keep up and the pair starts to diverge.
Figure 1: The solid curve shows the characteristic timescale for the evolution of $a_{R}$ as a function of time. We see that this is initially much longer than the lifetime of the disk, but shortens as time increases, i.e. the torque reversal moves out faster as $\dot{M}$ drops. The dotted curve represents $\tau_{a}$ at a location $0.63a_{R}$, assuming a mass $6M_{\oplus}$. This represents the rate of outward migration for a planet of that mass, trapped in a 2:1 resonance with a planet located at the torque reversal location. The dashed line represents the same but for a $3M_{\oplus}$ planet located at $0.763a_{R}$, i.e. in a 3:2 resonance. The long dashed line indicates the timescale for inward migration, evaluated at $a_{R}$. Both timescales are shorter than $\tau_{R}$ until about 1.2 Myr. After this, the planets find it harder to track $a_{R}$ and will freeze out and decouple from the disk.
The dashed line shows the case of a $3M_{\oplus}$ planet locked in a 3:2 resonance with an outer member located at $a_{R}$. This shows the same characteristic behaviour, although the divergence starts earlier.
The specific value of $a_{R}$ also depends on the stellar magnetic field and the degree to which it diffuses into the disk. For a simple dipole field (n=3), equation (4) is equivalent to a magnetic field on the stellar surface of 300 G. The magnetic fields observed in T Tauri stars can reach several kG (Johns-Krull, 2007) but many stars also show smaller or unobservable fields (Donati et al., 2011; Villebrun et al., 2019). Furthermore, Yu et al. (2023) showed that the results of the disk simulations scaled with the enclosed magnetic flux within the planetary radius. Thus, equation (4) also represents a surface field of $B_{0}=1.7$ kG with an effective n=3.5. Given this ambiguity in trading off $B_{0}$ and $n$, and the range of potential natal stellar fields, we shall investigate the effects of varying magnetic field by simply scaling the value of $a_{R}$ in subsequent sections.
A full description of the interaction between the star and the disk must also account for the fact that there is a true inner boundary at the edge of the magnetospheric cavity. In most of the cases we discuss here, the planets do not penetrate in far enough to reach this boundary, and so we will simply consider the magnetospheric cavity to lie at $a_{R}/3.7$ unless otherwise noted.
3 Converging versus Diverging
When two planets converge during migration, they enter a state of resonant lock, in which the rate of change of orbital frequencies are equal, i.e. $\dot{n_{1}}/n_{1}=\dot{n_{2}}/n_{2}$. If we simultaneously impose the condition that the libration amplitude of the resonant angle is small, the eccentricities reach equilibrium values given by (Terquem & Papaloizou, 2019), in the ratio
$\left(\frac{e_{2}}{e_{1}}\right)^{2}=\frac{1}{4}\left(\frac{m_{1}}{\alpha_{12}m_{2}}\right)^{2}\left(\frac{f_{2}^{\prime\prime}}{f_{1}}\right)^{2}$ (5)
where we have chosen the q=1 resonance for this example and the various quantities are defined in appendix A. These conditions also imply specific values for the equilibrium eccentricities. The full expressions are given in Terquem & Papaloizou (2019), but the most important element is that
$e_{1}^{2}\propto\left(\frac{t_{e,1}}{t_{a,2}}-\frac{t_{e,1}}{t_{a,1}}\right)$ (6)
In the traditional case where both planets are migrating inwards, this quantity is positive as long as $t_{a,1}>t_{a,2}$, i.e. convergent migration yields resonant lock. In our model, the torques can reverse sign, which can be modelled as simply reversing the sign of the relevant timescale.
Thus, when the innermost planet first crosses the torque reversal location, the value of $t_{a,1}$ will change sign. This will cause the pair to converge even more strongly, so that there will still be an equilibrium, although the values of the equilibrium eccentricities will change. However, once the second planet crosses the torque reversal location, then both planets are moving outwards and the question of whether they are converging or diverging is once again a matter of their relative motion. The equilibrium will fail if the quantity in equation (6) becomes negative – thereby breaking the resonant lock. The system evolution retains the same character if one of the planets is tracking the location of the torque reversal – the relative $t_{a}$ is then simply dictated by the evolution of the torque reversal.
Extending the analysis to three planet systems yields qualitatively similar behavior, with locked resonant systems exhibiting equilibrium eccentricities in fixed proportions, but also demonstrating failure of these equilibria once the planets start to move outwards again.
4 Two Planet Systems
Figure 2 illustrates the basic evolution for a two planet system, containing two $3M_{\oplus}$ planets. The black curve in the upper left plot illustrates the evolution of the semi-major axis of the inner planet, and the red curve the evolution of the outer planet. The dotted line indicates the evolution of the location of the torque reversal location $a_{R}$. The upper right panel shows the period ratio of the two planets as a function of time. In the lower left, we show the evolution of the planetary eccentricities, and, in the lower right, the panel shows how the resonant angle for the inner and outer 3:2 resonances change with the planetary and disk evolution.
Figure 2: The upper left panel shows the evolution of semi-major axis for a pair of $3M_{\oplus}$ planets migrating due to the gaseous torque model described in the text. The dotted curve shows how the location of the torque reversal evolves with time. Important transitions in the dynamical state of the system are labelled as A–E, and are show as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we see the evolution of the resonant angles $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-\bar{\omega}_{2}$ (red).
The first incident of note – denoted point A – is when the inner planet reaches the torque reversal location $a_{R}$. At this point the inward migration of the inner planet stops and the planet semi-major axis starts to track the evolution of $a_{R}$, which is slowly moving outwards. Up until this point, the period ratio is increasing (the planets are diverging) because the migration time decreases as planets get closer to the star. However, with the halting of the inner planet migration at A, the planets begin to converge because the outer planet, shown as the red curve, continues to migrate inwards, and the two planets enter the 3:2 resonance at B. The net relative motion is convergent and so the pair become locked in a resonant configuration. They also acquire finite equilibrium eccentricities, as expected from the discussion in § 3.
The torques on the outer planet now drive the pair inwards, and the inner planet is pushed interior to the torque reversal. Eventually the outer planet reaches the torque reversal (at time denoted as C). The planets remain in equilibrium (because the torques on the inner planet are pushing it outwards, so the pair is still converging) but the equiliibrium eccentricities change because the sign of the torques change. After this, the pair starts to migrate outwards but remains in resonance as long as the rate of outward migration of the inner planet is faster than the evolution of the torque reversal. This continues for a long time, but eventually, as the torques weaken and the outward motion of $a_{R}$ accelerates, the pair starts to diverge, at epoch D. The libration in the 3:2 resonance begins to increase, as does the period ratio, and the eccentricities start to damp away as their resonant excitation weakens. Eventually the migration freezes out completely (at E), leaving the system still in resonance with a finite libration amplitude and an offset relative to strict commensurability.
This is the fundamental result that emerges from these calculations – in the disk models of Yu et al. (2023), planets are not trapped at the true inner edge of the disk, but rather at the location of the torque reversal. As the disk evolves and the surface density drops, this location moves outwards, resulting in an outward migration of the planets at, and interior to, the torque reversal. In many cases, this leads to divergent migration, thereby breaking the resonant lock and causing divergence of the period ratios away from the initial commensurabilities, to values more consistent with the observations.
The planetary masses also play a role in determining the outcome of the evolution, because more massive planets remain coupled to the gas disk for longer, and lower mass planets freeze out earlier. To illustrate this, Figure 3 shows the case of a two planet system wherein the outer planet ($6M_{\oplus}$) is more massive than the inner one ($3M_{\oplus}$). The initial evolution is qualitatively similar, except that the stronger torques on the outer planet means that the inner planet barely halts at the torque reversal before being captured in resonance and pushed further inwards. The eventual divergence from the 3:2 resonance (at D) is also qualitatively similar. However, the outer planet remains coupled to the gas disk for longer by virtue of its larger mass and so migrates far enough that the system passes through the 2:1 resonance (shown at D). This resonant encounter is now divergent, so there is no capture (the system leaves the 2:1 resonance at E) but there is a transient excitation of the eccentricities that occurs late enough to leave small non-zero eccentriities ($\sim 0.01$) after the system has frozen out (at F).
Figure 3: The upper left panel shows the evolution of semi-major axis for a pair of migrating planets where the outer planet is $6M_{\oplus}$ and the inner planet is $3M_{\oplus}$. The dotted curve shows how $a_{R}$ evolves with time. Important transitions in the dynamical state of the system are labelled as A–F, and are shown as vertical dashed lines in the other three panels. In the upper right, we see the evolution of the period ratio for the pair. In the lower left we show the evolution of the eccentricity for each planet (red curve indicating the outer planet and black curve the inner planet). In the bottom right, we have two panels to track the evolution of the resonant angles for the two important first order resonances that are encountered. The one designated as $\phi_{3:2}$ shows the angles $3\lambda_{2}-2\lambda_{1}-2\bar{\omega}_{1}$ (black) and $3\lambda_{2}-2\lambda_{1}-2\bar{\omega}_{2}$ (red). The one designated as $\phi_{2:1}$ shows the evolution of the two resonant angles $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{1}$ (black) and $2\lambda_{2}-\lambda_{1}-\bar{\omega}_{2}$ (red).
The overall amount of migration is driven largely by the planetary masses, as they determine how long the planet remains coupled to the disk and how rapidly the planets migrate with respect to one another. Figure 4 shows the effect of both planetary mass ratio and total mass in the overall evolution of a range of planet pairs. We see that low mass ratios ($M_{2}/M_{1}<0.6$) leave pairs trapped in resonance, whereas larger mass ratios lead to divergence and a wide range of final period ratios. This is a consequence of the fact that the more massive planets remain coupled to the gas for longer. Thus, if the inner planet is the more massive one, the pair will remain in the convergent state for the entirety of the disk evolution. However, if the outer planet is the more massive one, then the inner planet will freeze out first and fall behind the outer planet’s migration. This leads to a diverging configuration and a final period ratio that increases as $M_{2}/M_{1}$ does. The degree of divergence also depends on the total planet mass, with more massive planetary systems leading to large final mass ratios (because they remain coupled to the disk for longer).
Figure 4: The lower panel shows the final period ratio of a planetary pair as a function of the mass ratio $M_{2}/M_{1}$. The results are color coded according to the final value, with black signifying $P_{2}/P_{1}<1.55$, red (1.55–1.95), magenta (1.95–2.05), green (2.05–3) and cyan for $P_{2}/P_{1}>3$. In the upper panel, these colors indicate the final states of systems as a function of mass ratio and total system mass $M_{1}+M_{2}$. The dotted lines in the lower panel indicate the first order commensurabilities (1+q)/q for q=2 and q=1.
4.1 Dependance on model parameters
The evolution of these planetary pairs is driven by the torques due to the gas disk, so they will be influenced by the parameters of the gaseous disk, and also by the strength of the stellar magnetic field – which determines the location of the torque reversal that drives the outward motion and divergence. Figure 5 illustrates this.
Figure 5 shows the evolution of a pair of $6M_{\oplus}$ planets in three different scenarios. In each case, the accretion disk evolution is the same, but we have scaled the form of $a_{R}$, given in equation (4), by a constant factor. As discussed above, this mimics variations in either the stellar magnetic field strength, or the degree of field compression by the inflow ($n$). Thus the black curve in Figure 5 shows the default case, while the blue curve shows the case where $a_{T}$ is a factor 0.7 smaller (corresponding to a weaker magnetospheric influence) and the red case represents the case where $a_{T}$ is a factor 1.3 larger (corresponding to a stronger magnetospheric influence).
Figure 5: The upper curves show the evolution of the semi-major axes of pairs of $6M_{\oplus}$ planets evolving in disks with different torque reversal locations ($a_{R}$). The black curve corresponds to the model given by equation (4). The red curves correspond to $a_{R}$ scaled up by a factor 1.3, and the blue curves have $a_{R}$ scaled down by 0.7. The dotted lines show the evolution of the respective $a_{R}$. The lower panel shows the evolution of the corresponding period ratios.
We see that more distant torque reversals result in an earlier entry into resonance, and a slightly later departure from commensurability. The end result is a more widely separated pair of planets. Given that we expect a range of stellar magnetic fields amongst exoplanet host stars, this will also result in some variation of the planetary spacings.
The final separations also depend on the model for the evolution of the accretion rate and surface density of the disk. Our default model assumes an evolution characterised by two exponentials, chosen to mimic the slow decay of a viscously evolving disk that accelerates once the inner disk has been decoupled from the outer disk by photoevaporation. Figure 6 shows how the evolution of a pair of $6M_{\oplus}$ planets changes depending on the choice of the faster exponential cutoff. We show the cases where the cutoff rolls after after 1.7 Myr, 1.9 Myr or not at all (i.e. the decay is characterised by only one exponential – the slower one). We see that later cutoffs lead to more divergence.
Figure 6: The curves show the evolution of a pair of $6M_{\oplus}$ planets in the same disk, with the same initial conditions, but with three different values for the time ($t_{c}$) at which the accretion rate starts to rapidly drop off. In the upper left panel, the three black curves represent the inner planet and the three red curves the outer planet. The upper right hand panel shows the evolution of the respective period ratios. In the bottom left, the three panels show the evolution of the resonant angles for the 3:2 resonance (with $t_{c}=1.7$Myr on top, $t_{c}=1.9$ Myr below that and $t_{c}=\infty$ on the bottom). The lower right panels show the corresponding evolution for the 2:1 resonant angles.
5 Three Planet Systems
In the previous section, we have demonstrated that the presence of a transition region and torque reversal in the inner disk can cause a significant dynamical shake-up in planetary systems, potentially explaining the low frequency of observed resonant pairs. The sample of planets discovered by the Kepler satellite contains many systems with multiplicity higher than two. In this section we will examine the effects of a torque reversal on three planet systems.
The addition of a third planet adds some complexity to the system, as the planet in the middle now experiences gravitational interactions from both the inner and outer companions. As a result, this planet may be evolving outwards with the torque reversal – and diverging from the inner planet – but then experiences torques from the third, outer, planet, which is still being driven inwards. This influence can counter the evolution discussed in the previous section, and so we must consider the parameters associated with this competition.
Figure 7 shows the simple extension of our model system shown in Figure 2, in which we simply add a third planet, also of mass $3M_{\oplus}$. The initial configuration is chosen so that both pairs are between the 3:2 and 2:1 commensurabilities. The initial evolution is divergent, because the rate of migration is faster closer to the star. The first planet encounters the torque reversal at A, and so the inner pair begins to converge, trapping into the 3:2 resonance at B. This excites the eccentricities of the inner pair to equilibrium values, but the outer planet remains almost circular as it continues to migrate inwards. The resonant interaction slows the inward migration of the middle planet and the outer pair encounters the 3:2 resonance at C. This also briefly disturbs the resonance lock of the inner pair, but both pairs soon settle into resonance again when the middle planet reaches $a_{R}$ (at D). The outer planet continues to migrate inwards, pushing both the inner and middle planets interior to $a_{R}$, until it eventually encounters the torque reversal as well (at E). At this point we have three planets in a resonant chain of 3:2 resonances, moving outwards with the outer planet tracking $a_{R}$. As we have noted above, this will continue until the evolution of $a_{R}$ accelerates and the planets start to diverge (at F). The divergence is not rapid, and the planets eventually freeze out (at G) still in resonance, with a small libration, and moderate separations from commensurability.
Figure 7: Here we show a three planet system, containing three planets, each with mass $3M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the resonant angles for the 3:2 resonance for both the outer pair (red and blue) and inner pair (red and black). The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
The end result is a trio of planets, with the inner pair having period ratio $P_{2}/P_{1}=1.54$, and the outer pair with a period ratio $P_{3}/P_{2}=1.57$ – similar to what one would expect by simply combining the two pairs independently.
Figure 8 shows the effect of scaling all the masses up by a factor of 2 – i.e. a system containing three planets of mass $6M_{\oplus}$. The inner pair evolves in the same manner as before – converging after the inner planet reaches the torque reversal (A), locking into the 3:2 resonance (B) and finding a new equilibrium when the middle planet reaches the torque reversal (C). However, the stronger gravitational interactions result in the outer pair locking into the 2:1 resonance, at D. Furthermore, the interactions are strong enough to prevent the outer planet from forcing its way inward to $a_{R}$. Thus, the system evolves for a period of time with the middle planet tracking $a_{R}$ and maintaining an inner 3:2 and outer 2:1 resonance.
Figure 8: Here we show a three planet system, containing three planets, each with mass $6M_{\oplus}$. The evolution of the inner planet is shown in black, the middle planet in red, and the outer planet in blue. In the upper right panel, the red curve refers to period ratio of the inner pair, while the blue curve indicates the period ratio of the outer pair. In the lower right, we show the behaviour of the resonant angles for the different first order resonances at play. The upper pair of panels in this quadrant show the evolution of the 3:2 resonance for both the outer pair (red and blue) and inner pair (red and black). The bottom pair of panels show the corresponding evolution for the 2:1 resonance. The labels in the upper left indicate important epochs in the evolution of the system, identified as vertical dashed lines in the other panels.
Eventually, as the outwards evolution of $a_{R}$ accelerates, the middle planet detaches from the torque reversal and the outer planet resumes its inward migration (E), eventually reaching $a_{R}$ at F. At this point the evolution is similar again, with the planets starting to diverge from strict commensurability as the outer planet decouples from the torque reversal (G) and eventually freezing out at H. One noteworthy difference here though, is the behaviour of the outer pair, wherein the outer 2:1 resonant angle reverses its libration orientation. This occurs when the period ratio diverges beyond $P_{3}/P_{2}=2.25$, at which point the stable libration equilibrium switches in angle by $\pi$ (as discussed in appendix A).
5.1 Observational Comparison
These calculations demonstrate that the evolution of systems of multple planets in the protoplanetary disk background calculated by Yu et al. (2023) can produce divergence from resonance and result in final period ratios well away from the commensurabilities expected in a naive application of the resonant locking theory.
In a perfect scenario, we should be able to model the divergence of specific planetary systems with well-defined masses, although the stellar magnetic field value and the evolution of the protoplanetary disk remain quantities that need to be modelled and so can introduce some variations. Of course, the observational reality is less well constrained – most of the known multiplanet systems are detected via the transit method, and most of the planetary masses are small enough to prove challenging to measure with the radial velocity method,
Nevertheless, it is instructive to compare our existing model with the few known planetary systems where the planetary masses are well measured. As a comparison for our model, we require systems of three planets, orbiting approximately Solar-mass stars, and with well measured masses. We restrict ourselves solely to triple planet systems, since the presence of neighbors does influence the degree of divergence.
Figure 9 shows a comparison for the two planetary systems that best match our criteria. In the top panel we show a model for the planetary system around the star HIP29442. This system has a host star that it is close to solar mass ($0.89M_{\odot}$), and has exactly three planets with well-constrained masses (errors $<1M_{\oplus}$). In the bottom panel, we show a comparison to the star GJ9827. The host star in this system is $0.62M_{\odot}$, which is probably too low to be strictly applicable to our model, but is the next closest match and still instructive. Other systems with well measured masses either orbit M dwarfs or have either more or less than three planets.
Figure 9: The curves in the upper two panels show the evolution of a planetary system modelled after the HIP29442 planetary system, assuming our standard disk and a torque reversal scaled by a factor 0.73. The horizontal dashed lines in the left hand panel indicate the observed final semi-major axes and the horizontal dashed lines in the upper right panel indicate the observed final period ratios. The lower panels show the equivalent evolutions for the GJ9827 system, in which the torque reversal is scaled by 0.55. In both panels, the upper dotted line indicates the torque reversal $a_{R}$, while the lower dotted line indicates the edge of the magnetospheric cavity, where the torques go to zero.
The planetary masses for HIP29442 measured by Damasso et al. (2023) are $4.5\pm 0.3M_{\oplus}$, $5.1\pm 0.4M_{\oplus}$ and $9.6\pm 0.8M_{\oplus}$ for the three planets with orbital periods of 3.538 days (c), 6.429 days (d) and 13.63 days (b). To model this system we choose the scaling of $a_{R}$ so that the innermost planet freezes out at approximately the correct distance. The resulting evolution shows the system enters a 3:2+3:2 resonant chain early and then diverges to final period ratios that are similar to those observed, although slightly smaller. In particular, we note that the outer pair diverged from interior of the 2:1 resonance to outside it.
The masses for the GJ9827 planets (Passegger et al., 2024) are $4.3\pm 0.3M_{\oplus}$, $1.9\pm 0.4M_{\oplus}$ and $3.0\pm 0.6M_{\oplus}$, corresponding to orbital periods of 1.209 days (b), 3.648 days (c) and 6.202 days (d). In this case, we scaled $a_{R}$ so that the middle planet freezes out at the correct distance, and scaled the accretion rate down by a factor of 0.7 to account for the lower stellar mass. With these adjustments our migration model is able to match the outer pair quite well, but the inner planet does not migrate in as far as desired. We have also moved the edge of the magnetospheric cavity out to a value of $a_{R}/2$. This halts the migration of the inward planet once it crosses into the cavity, but the inward migration is still halted by the reversal of the outermost planet when it hits $a_{R}$. Thus, the inner pair mass ratio falls somewhat short of the observed value, but does reproduce the qualitative features of the system in that the inner pair is wider than the 2:1 resonance while the outer pair lies between 3:2 and 2:1.
Of course, most planetary systems do not have masses measured to this accuracy, but our principal goal is not to constrain the histories of individual planets, but to understand the features of the underlying population of planetary systems. Thus, we wish to take a closer look at the results of the Kepler mission, when interpreted in the context of our model.
5.2 Kepler Analog Systems
As we can see from Figure 4, the degree of divergence of a planetary pair away from resonance is partially regulated by the mass ratio. Thus, to survey a realistic range of masses, we choose to simulate all mass combinations detected amongst Kepler triple systems. The Kepler results provide radii, not masses, so we must use an empirical mass-radius relation to convert each set of radii to masses. We use the results of Chen & Kipping (2018), which provides mass estimates for 7000 KOI, based on the model outlined in Chen & Kipping (2017). In particular, this model includes a spread in the expected masses at fixed radius, in accordance with the observational evidence. However, the possible spread in masses from planet to planet is quite large, and selecting three masses independently will give a large variation in mass ratios – inconsistent with the empirical peas-in-a-pod phenomenon (Weiss et al., 2018), which states that the degree of uniformity in planetary properties within a given system is smaller than that within the overall exoplanet population as a whole.
As a result, we choose the mass of the middle planet from the results of Chen & Kipping (2018) and then scale the inner and outer planets relative to this using the deterministic component of the mass-radius relation in Chen & Kipping (2017). This allows for some variation in the overall masses, but keeps the internal mass variation consistent with the radius variations. Using these masses we evolve forward realisations of the 107 Kepler planetary systems that contain three confirmed planets, none of which have a radius $>6R_{\oplus}$ (our torque model is not applicable to Jupiter-class planets, which will open gaps in the protoplanetary disk). We wish to focus on the effects of the outspiral and detachment from the disk at late times, so we assume the planets occupy an initially compact configuration, starting the planets at 0.076 AU, 0.107 AU and 0.158 AU, so that the systems lie between the 3:2 and 2:1 resonances.
The results of this parameter scan is shown in Figure 10. The crosses show the parameters of the known Kepler three planet systems – with inner pair period ratio on the horizontal axis and outer pair period ratio on the vertical axis. The solid red points show the results of the model calculations. We have divided this parameter space into four quadrants based on where the period ratios lie relative to the first order resonances (q+1:q). Systems in quadrant A have both pairs between the q=1 and q=2 first order resonances. Systems in quadrant B have a compact inner pair (between q=1 and q=2) and a wider outer pair (wider than q=1). Quadrant C has both pairs wider than the q=1 resonance and quadrant D has a wider inner pair (outside q=1) and a compact outer pair (between q=2 and q=1).
Figure 10: The solid red points represent the period ratios of inner (horizontal axis) and outer (vertical axis) pairs from our simulations of the set of Kepler triple analoges. The crosses correspond to observed Kepler triples, and the dotted lines indicate the 3:2 and 2:1 commensurabilities. The labelling of the four quadrants A–D are discussed in the text.
We see that our model is capable of producing systems that lie in all four quadrants, as well as a handful that cluster near the resonant chains. This variation is achieved solely as a function of the different masses – all of the red points are generated for the same disk evolution model, initial spatial configurations and $a_{R}$ history. This demonstrates that the mass variations in the Kepler systems are capable of producing a substantial fraction of the overall period ratio variations.
Figure 11 shows how the mass ratios influence the outcome of these models. It shows the distribution of mass ratios for model planetary systems in the four different quadrants. The fact that the systems in Figure 7 and Figure 8 evolve to quadrants A and B respectively, despite starting at the same point (the intersection of the dotted lines in Figure 11), illustrates that total mass plays a role as well, but Figure 11 shows that the mass ratios are an important factor.
Figure 11: The four panels show the planetary mass ratios for model three planet systems that are found in each of the four quadrants identified in Figure 10. Those pairs denoted as solid points are still located close to a commensurability (within 1.50–1.55 or 2.00–2.05) at the end of the simulation, while those represented as open circles are well separated. The dotted lines indicate equal mass ratios for the inner pair (vertical) and outer pair (horizontal). text.
Objects in quadrant A tend to have the outermost planet larger than the middle planet, and those that are found far from resonance also tend to have the middle planet larger than the inner one. The planets in quadrant B tend to be found with middle planets larger than inner planets. Those planets in quadrant C tend to exhibit an overall trend of increasing mass outwards, while those in quadrant D tend to have inner pairs with similar mass ratios.
We have already seen examples of evolutionary tracks for quadrants A and B in Figures 7 and 8. Figure 12 shows the example of one that ends up in quadrant C. This is a relatively massive system, with planets of mass 8.1, 9.7 and 10.5 $M_{\oplus}$, a realisation of the masses in the system Kepler-60. It shows that the rapid inward migration causes the inner pair to diverge to the point that they are wider than the 2:1 resonance. When the inner planet halts (at A) this trend reverses and the inner pair lock into the 2:1 resonance (at B). This also slows the inward migration of the middle planet so that the outer pair start to converge as well. However, this happens just before they diverge beyond 2:1, so that the convergence ultimately leads to their trapping in the 3:2 resonance. This transition leads to a a period of dynamical rearrangement, as the inner pair loses its resonant lock at C, before the outer pair actually locks into the 3:2 (at D) and only appears to settle down again once the outer planet reaches the torque reversal (at E).
Figure 12: In this plot, $M_{1}=8.1M_{\oplus}$, $M_{2}=9.7M_{\oplus}$ and $M_{3}=10.4M_{\oplus}$. The colours identify the planets in the same way as in Figure 7. We see that this system starts in the same configuration as the system in Figure 7, but the outer planet diverges so far that it crosses the 2:1 resonance, leaving both pairs wide of the 2:1 commensurability. As shown in the lower right panel, this capture is only temporary but does result in an eccentricity excitation (event G).
After this, the evolution is more stable, with a 2:1+3:2 chain evolving outwards until the system starts to diverge at F. The larger planetary masses means that the system takes longer to decouple, and the outer pair passes through the 2:1 resonance (at G). The system doesn’t trap because the evolution is divergent, but the resonant crossing does produce an eccentricity excitation that doesn’t entirely die away before the system completely freezes out (at H). The end result is that neither pair remains in resonance at the end.
Figure 13: This three planet system is for $M_{1}=4.7M_{\oplus}$, $M_{2}=4.1M_{\oplus}$ and $M_{3}=5.3M_{\oplus}$. The colours refer to the planets in the same order as in Figure 7. In this system, both planets capture into the 3:2 resonance and remain there as each of the planets reach the torque reversal. We see that this system remains in resonance at the end, even though the period ratios have diverged from 3:2 by then.
Figure 13 shows the evolution of a planetary system modelled after Kepler-295, with planetary masses of 4.7, 4.1 and 5.3 $M_{\oplus}$ respectively, representative of a system that ends up in quadrant D. In this case, the evolution is quite similar to that seen in Figure 12, with the inner pair captured into 2:1 after an initial divergence, and the outer pair captured into the 3:2, with a similar brief period of dynamical reshuffling. However, the outer planet is less massive in this case, and the system does not evolve far enough to pass through the 2:1 commensurability, although the outer pair does actually enter the 2:1 resonance by virtue of the large precession at low eccentricities.
5.2.1 Initial conditions
Up to this point, we have assumed that the planets either formed close in or arrived from further out at an early stage, so that their final states are determined by the outward migration at late times. However, if the planets spiral in at later times, they may never reach the same compact configurations as we have simulated above.
Thus, to test this assumption, we have evolved the same 107 Kepler analog systems forward, but now starting them at semi-major axes of 0.16 AU, 0.29 AU and 0.54 AU. This does indeed lead to wider final configurations, as many systems get trapped in the 2:1 resonance during their inspiral, so that the detachment from resonance starts from a wider configuration. Furthermore, systems with low mass outer planets may never experience resonance locking because the inner planets migrate inwards faster and the outer planet cannot catch up before the disk evaporates. This can leave an inner pair near the resonance accompanied by a very wide outer pair – a system that will most often be observed simply as a double.
Figure 14 illustrates the effect of late time inspiral by showing the evolution of a system with the same mass configuration as in Figure 13, with the new, more distant, initial conditions. The inner planet reaches the torque reversal relatively early (at A), and the middle planet catches up to it at B. The systems started wider in this case and so they capture into the 2:1 resonance at this point and the torques on the middle planet drive the pair inwards until the middle planet reaches the torque reversal at C. Eventually, the outer planet gets close enough to encounter the 2:1 resonance with the middle planet at D.
Figure 14: This three planet system is for $M_{1}=4.7M_{\oplus}$, $M_{2}=4.1M_{\oplus}$ and $M_{3}=5.3M_{\oplus}$. The resonant angles at the lower right are for the 2:1 resonance for the inner pair (upper panel) and outer pair (lower panel). The system gets trapped into a 2:1+2:1 resonant chain that eventually diverges to a final position in Quadrant C.
This causes some dynamical shake-up in the inner resonance but there is a period of time where both pairs librate in a 2:1 resonance, albeit with non-zero amplitudes and changes in configuration (E). Eventually the outer planet reaches the torque reversal (at F). This occurs late enough that the pairs immediately start diverging. Although both pairs pass through period ratio 2.25 (at G), they are already out of resonance at this stage, and end up freezing out at H. At this point the system is well located in quadrant C.
6 Discussion
The results of § 5 demonstrate that our model results in the disruption of the majority of the resonant chains that form initially, in keeping with the qualitative features of the observed Kepler sample. To make a quantitative comparison, we must now compare the results of our simulations to the observations.
The Kepler satellite detects planets via their transits, so we must calculate the geometric probability that each of the planets in our simulations are actually observed. The probability of observing the transit of an individual planet is straightforward, scaling as the inverse of the semi-major axis. The relative frequency of transit of members of multiple planet systems depends on the inclination dispersion within the system. To assess these probabilities, we observe each system from $10^{5}$ random orientations on the sky. For each realisation, the mutual inclinations are drawn from a Rayleigh distribution with an inclination dispersion of $3^{\circ}$, the upper limit consistent with the Kepler statistics (Fang & Margot, 2012). The relative lines of nodes are assumed to be randomly distributed. With these results, we determine the relative frequency of observing all three planets in each system, and the frequency of observing the three different permutations of planetary pairs.
Figure 15 shows the resulting probability of observing different period ratios between neighbouring planets. In the upper panel, we show the observed distribution of Kepler period ratios. We show two samples. The first is the standard compilation, from Fabrycky et al. (2014), shown as a dotted histogram. Since this is now ten years old, and other systems have been confirmed since, we also show, as a solid histogram, those planets identified in the NASA Exoplanet Archive (Akeson et al., 2013) as confirmed, and discovered by Kepler. The two first order resonances (q+1)/q, with q=1 and q=2, are shown as dashed lines, illustrating the well-known result that most nearest neighbour pairs in the Kepler sample are not near a commensurability.
Figure 15: The upper panel shows the observed distribution of nearest neighbour period ratios in the sample of Kepler systems, based on the compilation of Fabrycky et al. (2014) (dotted histogram) and the latest compilation of confirmed planets, discovered by Kepler, from the NASA exoplanet archive. The three panels below that show the outputs of our simulations. The second panel shows the final distributions that started with compact initial conditions and mass ratios drawn from the observed Kepler triple systems, and assuming that all stars the default value of the transition radius $a_{R}$. The third panel shows the results of similar initial conditions, but adding together systems drawn equally from $a_{R}=0.8$, $a_{R}=1.2$ and $a_{R}=1.5$. The bottom panel shows the results from the simulations that begain with more distant, and wider spaced, initial conditions. This means that more systems get trapped into the 2:1 resonance instead of the 3:2 resonances, and the later divergence leads to wider systems. These simulations assume the standard $a_{R}$.
The other panels in Figure 15 show the output from our simulations. The second panel from the top, labelled as ‘Compact’, represents the results from our default simulations (the ones shown as red points in Figure 10). These have been assigned weights according to their geometric probability of observation, as discussed above. The third panel from the top represents the evolution of the same set of initial conditions, but with $a_{R}$ scaled up by a factor of 1.2 relative to the default value given in equation (4). This represents the case of a stronger stellar magnetic fields. The resulting distribution is pretty similar, although the resonant peaks are less pronounced – i.e. a larger fraction of systems moved away from commensurabilities.
Finally, the lower panel shows the results of starting with wide initial conditions (and an $a_{R}$ corresponding to that of the second panel, i.e. all systems have the default $a_{R}$). We see that many more of these systems get captured in the 2:1 resonance and we get a distribution that skews to wider separations than expected in the observations.
The most important feature of this figure is that the simulations produce a similar fraction of systems wide of the commensurabilities as seen in the upper panel. The compact configurations produce a lot of systems between the q=1 and q=2 commensurabilities, while the more distant initial conditions primarily produce systems wide of the q=1 commensurability.
A more quantitative comparison of the non-resonant fractions is shown in Figure 16, which compares the cumulative distribution of period ratios between theory and observations.
Figure 16: The black histogram shows the cumulative distribution of period ratios for the sample of current confirmed Kepler planetary systems – the summed version of the solid histogram in the upper panel of Figure 15. The red histogram shows the cumulative distribution from our default models – those shown in the second panel of Figure 15. The magenta histogram shows the models from the third panel of Figure 15, demonstrating that the simulations with larger $a_{R}$ produce too many systems in compact configurations. The blue histogram represents the lowest panel in Figure 15 – those derived from the simulations that start with wide initial conditions. The cyan histogram represents a sample composed of 70% from the magenta histogram and 30% from the blue sample, indicating that this too can produce a reasonably accurate representation of the observations.
The black histogram represents the observed distribution drawn from the current Kepler sample (the solid histogram in the upper panel Figure 15). The red histogram represents the default model sample from the second panel, and which reproduces the observed distribution quite accurately. The magenta histogram shows the distribution from the third panel. This one shows too many at lower period ratios and too few at higher period ratios. The blue histogram shows the sample from the lowest panel – those drawn from wide initial conditions. In this case the model clearly produces too few close pairs and even falls below the observed curve at larger periods, indicating a substantial pile-up of systems between period ratios of 2.2 and 3. Overall, we find that the default model does an excellent job of providing a distribution of period ratios broadly consistent with those observed. A Kolmogorov-Smirnov test indicates that the two distributions have a 70% chance of being drawn from the same underlying distribution. The observations can also be reproduced with combinations of the different samples. The cyan histogram shows a model in which 70% of the systems come from the simulations with larger $a_{R}$ and compact initial conditions, along with 30% of the systems that began with wider initial conditions.
It is instructive to compare these results with those of prior related studies. The magnetospheric rebound model of Liu et al. (2017) describes a model of divergent migration driven by a one-sided torque exerted at the disk inner edge truncation, which can drive the outer planet away from an inner companion and break the resonance. Simulations of the Kepler population using this model (Liu & Ormel, 2017) does indeed produce a population of non-resonant planetary pairs, but still leaves too many planets in resonant lock. Our model allows for multiple planets to experience outwardly directed migration, and the divergence arises when the relative rates of migration start to differ. Comparison of our simulation results with those of Liu & Ormel (2017) shows that our model produces more widely spaced systems and a much better agreement with observation, as shown in Figure 16.
Another useful comparison set is the simulations of Izidoro et al. (2017) and Izidoro et al. (2021), the most comprehensive effort to date to compare the migration model with observations in a quantitative fashion. Izidoro et al. (2017) describe a model for the migration of low mass planets in both static and turbulent disks, which decay away on Myr timescales, leaving planets trapped in a chain of mean motion resonances. They furthermore run the simulations forward a further 100 Myr after the dispersal of the gas disk, and track the evolution of systems that undergo dynamical instabilities which can break the resonant chains. The encouraging part of the results of Izidoro et al. (2017) is that the resulting instabilities do indeed produce many non-resonant systems. However, the fraction of resonant systems that remain is still far too high ($\sim 50\%$) and the dynamical instabilties lead to planetary collisions which result in systems too massive compared to the observed systems. The simulations of Izidoro et al. (2021) takes these models further by including a prescription for gas-assisted pebble accretion during the migration phase. Variation in the pebble flux and the formation time of planetesimals result in various levels of dynamical instability. The authors find that it is possible to match the neighbour period ratio distribution of the Kepler planets but, once again, only if one treats the fraction of unstable systems as a free parameter. For some parameter choices, most of the simulated systems do go unstable, but these also, once again, produce systems that are too massive.
Our results show that both of these problems can be alleviated with a more physically motivated inner boundary condition. We note that our approximate match to the observations in Figure 16 uses the full set of simulated systems, so that our remaining resonant fraction is consistent with observations. Furthermore, these results are achieved using planetary masses estimated directly from the observations, and so demonstrate that the period ratios and masses can be reproduced in a mutually consistent manner. This approach does leave open the question of how and where these planets assembled, but it demonstrates that their final orbital configurations are largely determined by the manner in which the disk evolves and dissipates.
The high multiplicity of these compact planetary systems has spurred many discussions regarding their dynamical stability. The question of whether the planetary systems are on the edge of dynamical instability is often expressed in terms of the planetary spacing as measured in terms of mutual Hill radii. It has been claimed that the Kepler multiplanet systems are close to being ‘maximally packed’ (Lovis et al., 2011; Fang & Margot, 2013; Pu & Wu, 2015), in the sense that they cannot be packed significantly closer without being dynamically unstable. Figure 17 shows the final spacings of our model three planet systems in terms of mutual radii. We see the final spacings of these planetary triples peaks between 10 and 20 Hill radii, which is similar to that estimated for Kepler systems.
[…]
Indeed, the distribution is actually quite similar to that which emerges from simulations that model the planetary configurations emerging from in situ assembly. The dotted histogram shows the outcomes from the simulations of Hansen & Murray (2013) for the case of a total solid inventory of $20M_{\oplus}$. The magnetospheric rebound model actually produces systems that are even more compact than the in situ models, despite the fact that the latter are partially sculpted by considerations of dynamical stability, while the spacing of the former are regulated by the freeze-out process of the gaseous disk.
The time at which a given planet freezes out of the migration will also determine the final orbital period of the planet. This will also be affected by the value of $a_{R}$ and hence the stellar magnetic field. To test whether our model values are reasonable, we smulated the same 107 Kepler systems with the default value of $a_{R}$, a version with $a_{R}$ scaled up by a factor 1.2 (discussed above already) and also one with a scaling factor of 0.8. Figure 18 shows the resulting distribution of final planetary locations for all planets summed over these three simulations. In this case we do not adjust the counts for the geometric probabilities because the observations to which we compare have already accounted for selection effects and purport to represent the true underlying distribution. The shaded regions are these corresponding observed trends in planet occurrence for Super-Earths, as measured by Petigura et al. (2018). The simulations provide an excellent match to the observed rise from short periods to a peak in the range of $\sim$ 6–8 days, followed by a drop-off towards larger periods. Immediately exterior to the peak, the simulations also match the slope measured by Petigura et al. (2018) but begin to drop below this at orbital periods $>$20 days. This is a reflection of our restriction to triple systems. Systems with additional planets are expected to further fill in the distribution at longer periods.
Figure 18: The solid points indicate the number of planets in our simulations that end their evolution at a given orbital period. The error bars are Poisson counting errors. The shade regions indicate the observationally measured planetary occurrence rates for super-Earth planets from Petigura et al. (2018). The shaded regions indicate the range of slopes that lie within the observational uncertainties and the shaded regions have been scaled to match at the peak.
It is particularly noteworthy that the slope of the rise is matched well by these simulations. Previous attempts to explain this power law behaviour include connecting it to a range in stellar rotation rates (Lee & Chiang, 2017) via a magnetically or centrifugally maintained disk inner edge (Koenigl, 1991; Kenyon et al., 1996; Eisner et al., 2005). In the case of these calculations, it is determined by the mass distribution of the underlying planetary systems, which will determine the degree to which a planet is forced interior to the torque reversal and then when it freezes out.
7 Conclusions
The results presented here demonstrate that a detailed treatment of the inner boundary of protoplanetary disks Yu et al. (2023) has a strong influence on the final configurations of planetary systems that migrate through this disk. In particular, these models show the existence of an inner transition region of outwardly directed migration, which causes divergent migration of planets in resonant chains at late times. This results in a substantial fraction of initially resonant chains diverging to the point that their period ratios no longer lie close to first order commensurabilities (although some may still be technically resonant because of their high precession rates). The resulting distribution of planetary neighbour period ratios is a good match to the observed distribution of planets detected by Kepler, and also provides a good match to the observed relative occurrence frequency of observed Super-Earth planets. These results provide an explanation for the long-standing problem that the observed distribution of period ratios showed far fewer resonant systems than expected in the models.
Our model is not as complete as some of the models that preceded it (Izidoro et al., 2017, 2021), in that we do not present a full description of planetary evolution and assembly from planetary embryo formation up to final planetary architecture. Rather, our model is designed to emphasize that some of the most important features of the observations (such as planetary orbital periods and neighbour period ratios) are likely to be the consequence of the structure of the inner protoplanetary disk, and how the planets evolve as the disk disperses. This is encouraging in the sense that it implies that the result is robust with respect to some of the poorly understood physics involved in the assembly and early evolution of planetary systems.
On a less optimistic note, the fact that the period ratio distribution does not depend sensitively on the assembly process reduces the ways in which we can probe the formation conditions of planetary systems. It should be noted that our results match the observations if they begin with relatively compact configurations, so that examining the pathways to form the initial configuration may still provide a fruitful avenue of investigation. We will address this issue in a future publication.
Data availability: The data underlying this article will be shared on reasonable request to the corresponding author.
This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.
Appendix A A: Dynamical Equations
We wish to describe a system of three planets, subject to gaseous disk torques that cause migration in semi-major axis, and damping of eccentricities. The planets also interact gravitationally with their nearest neighbours, and our equations include the interactions with the first order resonances associated with the 2:1 and 3:2 commensurablities.
The planetary dynamics will be described in terms of the orbital frequencies, $n_{1}$, $n_{2}$ and $n_{3}$, normalised to a fidicual value $n_{0}$, so that $x_{i}=n_{i}/n_{0}$. These correspond to semi-major axis $a_{1}<a_{2}<a_{3}$, so that $\alpha_{12}=a_{1}/a_{2}$ and $\alpha_{23}=a_{2}/a_{3}$. Eccentricities are denoted by $e_{i}$ and the arguments of periastron are designed as $\bar{\omega}_{i}$. We will incorporate effects of the first order resonances $(1+q)/q$ for $q=1$ and $q=2$, between nearest neighbours. The resonant angles are $\phi_{1}=2\lambda_{2}-\lambda_{1}$, $\phi_{2}=3\lambda_{2}-2\lambda_{1}$, $\phi_{3}=2\lambda_{3}-\lambda_{2}$ and $\phi_{4}=3\lambda_{3}-2\lambda_{2}$. We normalise the timescale to $t_{0}=M_{*}/(3n_{0}m_{2})$, so that $\dot{x}_{i}=dx_{i}/d\tau=t_{0}dx_{i}/dt$.
The resulting evolutionary equations (Murray & Dermott, 2000; Goldreich & Schlichting, 2014; Terquem & Papaloizou, 2019) for the $x_{i}$ are
$\dot{x}_{1}$ $=$ $-x_{1}^{2}\alpha_{12}\left[A_{12}\sin\phi_{1}-B_{12}\cos\phi_{1}\right]-2x_{1}^{2}\alpha_{12}\left[C_{12}\sin\phi_{2}-D_{12}\cos\phi_{2}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{1}^{2}\frac{x_{1}}{\tau_{e,1}}$ (7)
$\dot{x}_{2}$ $=$ $2x_{2}^{2}\frac{m_{1}}{m_{2}}\left[A_{12}\sin\phi_{1}-B_{12}\cos\phi_{1}\right]+3x_{2}^{2}\frac{m_{1}}{m_{2}}\left[C_{12}\sin\phi_{2}-D_{12}\cos\phi_{2}\right]+$ (8)
$-x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[A_{23}\sin\phi_{3}-B_{23}\cos\phi_{3}\right]-2x_{2}^{2}\alpha_{23}\frac{m_{3}}{m_{2}}\left[C_{23}\sin\phi_{4}-D_{23}\cos\phi_{4}\right]+\frac{3}{2}\frac{x_{1}}{\tau_{a,1}}+3e_{2}^{2}\frac{x_{2}}{\tau_{e,2}}$
$\dot{x}_{3}$ $=$ $2x_{3}^{2}\left[A_{23}\sin\phi_{3}-B_{23}\cos\phi_{3}\right]+3x_{3}^{2}\left[C_{23}\sin\phi_{4}-D_{23}\cos\phi_{4}\right]+\frac{3}{2}\frac{x_{3}}{\tau_{a,3}}+3e_{3}^{2}\frac{x_{3}}{\tau_{e,3}}$ (9)
and the coefficients $A$$D$ are functions of the eccentricity,
$A_{ij}$ $=$ $f_{1}(\alpha_{ij})k_{i}+f^{\prime}_{2}(\alpha_{ij})k_{j}$ (10)
$B_{ij}$ $=$ $f_{1}(\alpha_{ij})h_{i}+f^{\prime}_{2}(\alpha_{ij})h_{j}$ (11)
$C_{ij}$ $=$ $f_{3}(\alpha_{ij})k_{i}+f_{4}(\alpha_{ij})k_{j}$ (12)
$D_{ij}$ $=$ $f_{3}(\alpha_{ij})h_{i}+f_{4}(\alpha_{ij})h_{j}$ (13)
and the quantities $f_{i}$ are the functions of the Laplace coefficients $b_{i}^{(j)}(\alpha)$, related to the expansion of the orbit-averaged disturbing function in terms of the resonant arguments. These are given by
$f_{1}(\alpha)$ $=$ $-\frac{1}{2}\left(4+\alpha\frac{d}{d\alpha}\right)b_{1/2}^{(2)}(\alpha)$ (15)
$f_{2}(\alpha)$ $=$ $\frac{1}{2}\left(3+\alpha\frac{d}{d\alpha}\right)b_{1/2}^{(1)}(\alpha)$ (16)
$f_{3}(\alpha)$ $=$ $-\frac{1}{2}\left(6+\alpha\frac{d}{d\alpha}\right)b_{1/2}^{(3)}(\alpha)$ (17)
$f_{4}(\alpha)$ $=$ $\frac{1}{2}\left(5+\alpha\frac{d}{d\alpha}\right)b_{1/2}^{(2)}(\alpha)$ (18)
$f_{2}^{\prime}(\alpha)$ $=$ $f_{2}(\alpha)-2\alpha$ (19)
$f_{2}^{\prime\prime}(\alpha)$ $=$ $f_{2}(\alpha)-\frac{1}{2\alpha^{2}}$ (20)
The eccentricity damping in these equations can often lead to low eccentricities which, in turn, leads to high precession rates. As a result, some of the resonant angles can continue to librate even when the period ratios deviate from the exact commensurability. Our equations represent the expansions in terms of the resonant arguments for the instantaneous osculating value of $\alpha$, and so we describe the functions $f_{j}$ as functions of $\alpha$ and not simply a number at exact commensurability.
Figure 19: The solid curves indicate the functions $f_{2}(\alpha)$, $f_{2}^{\prime}=f_{2}-2\alpha$ and $f^{\prime\prime}_{2}=f_{2}-1/(2\alpha^{2})$. The vertical dashed line indicates the value of $\alpha$ at which $f^{\prime\prime}_{2}$ changes sign. The upper axis shows the corresponding period ratio, indicating that the zero occurs at $P_{2}/P_{1}=2.252$.
Another consequence of the low eccentricities is that it is prudent to replace $e_{i}$ and $\bar{\omega}_{i}$ with $h_{i}$ and $k_{i}$, defined as $h_{i}=e_{i}\sin\bar{\omega}_{i}$ and $k_{i}=e_{i}\cos\bar{\omega}_{i}$. The resulting evolutionary equations are
$\dot{h}_{1}$ $=$ $-\frac{h_{1}}{\tau_{e,1}}+\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\cos\phi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\cos\phi_{2}$ (21)
$\dot{k}_{1}$ $=$ $-\frac{k_{1}}{\tau_{e,1}}-\frac{\alpha_{12}}{3}f_{1}(\alpha_{12})x_{1}\sin\phi_{1}+\frac{\alpha_{12}}{3}f_{3}(\alpha_{12})x_{1}\sin\phi_{2}$ (22)
$\dot{h}_{2}$ $=$ $-\frac{h_{2}}{\tau_{e,2}}+\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\cos\phi_{1}+\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\cos\phi_{2}$ (23)
$+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})\cos\phi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\cos\phi_{4}$
$\dot{k}_{2}$ $=$ $-\frac{k_{2}}{\tau_{e,2}}-\frac{1}{3}\frac{m_{1}}{m_{2}}f^{\prime\prime}_{2}(\alpha_{12})x_{2}\sin\phi_{1}-\frac{1}{3}\frac{m_{1}}{m_{2}}f_{4}(\alpha_{12})x_{2}\sin\phi_{2}$ (24)
$-\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{1}(\alpha_{23})x_{2}\sin\phi_{3}+\frac{\alpha_{23}}{3}\frac{m_{3}}{m_{2}}f_{3}(\alpha_{23})x_{2}\sin\phi_{4}$
$\dot{h}_{3}$ $=$ $-\frac{h_{3}}{\tau_{e,3}}+\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\cos\phi_{3}+\frac{1}{3}f_{4}(\alpha_{23})x_{2}\cos\phi_{4}$ (25)
$\dot{k}_{3}$ $=$ $-\frac{k_{3}}{\tau_{e,3}}-\frac{1}{3}f^{\prime\prime}_{2}(\alpha_{23})x_{3}\sin\phi_{3}-\frac{1}{3}f_{4}(\alpha_{23})x_{3}\sin\phi_{4}$ (26)
Finally, the above equations depend on the commensurabilities $\phi_{1}=2\lambda_{2}-\lambda_{1}$, $\phi_{2}=3\lambda_{2}-2\lambda_{1}$, $\phi_{3}=2\lambda_{3}-\lambda_{2}$ and $\phi_{4}=3\lambda_{3}-2\lambda_{2}$. The evolution of these variables is given by
$\dot{\phi}_{1}$ $=$ $\frac{1}{\Delta}\left(2x_{2}-x_{1}\right)$ (27)
$\dot{\phi}_{2}$ $=$ $\frac{1}{\Delta}\left(3x_{2}-2x_{1}\right)$ (28)
$\dot{\phi}_{3}$ $=$ $\frac{1}{\Delta}\left(2x_{3}-x_{2}\right)$ (29)
$\dot{\phi}_{4}$ $=$ $\frac{1}{\Delta}\left(3x_{3}-2x_{2}\right)$ (30)
where $\Delta=3m_{2}/M_{*}$.
One technical detail that results from our use of the full functional form of $f_{i}(\alpha)$ is that the function $f^{\prime\prime}_{2}$ changes sign if $\alpha<0.582$, which corresponds to a period ratio $>2.252$. This is shown in Figure 19 and is a consequence of the indirect contributions to the disturbing function. The sign of $f^{\prime\prime}_{2}$ determines the stable libration configuration of the 2:1 external resonance and so we see, in some of our results, a flip of the resonant libration as the period ratio increases. This only occurs if the resonance is well separated from the instability, which requires a fast libration and thus a small eccentricity. As a result, it is not clear this has an observable consequence but we include it for completeness.